How many g of Co(NO2)2 • 5 H2O are needed to make 432 mL of a 1.2 M solution?


1
Expert's answer
2021-06-02T06:04:34-0400

"M=1.2M"

"V=432ml"


"M=\dfrac{moles}{volume(ml)} \times1000"


"1.2= \dfrac{moles}{432} \times 1000"


"moles= 0.518"

Now:

"weight=moles \times molar" "mass"

"weight= 0.518 \times 210=108.78 gm"